OpenStudy (anonymous):

the population for a city is 65 438 and grows continuously at a rate of 4.7% each year. What is the approximate population in 22 years?

OpenStudy (anonymous):

I think you would just do 65,438x.47 get that answer. then, times that by 22...

OpenStudy (anonymous):

How ? are you 100% sure ?

OpenStudy (anonymous):

Oh no actually. there should be an equation but i don't know it. Becuase your compounding it. I just forgot the equation.. srry.

OpenStudy (anonymous):

Yeah, thats all wrong

OpenStudy (anonymous):

umm google like compounding interest equation and i think that would work.. i remember i did this last year..

OpenStudy (anonymous):

can you help me get the correct answer.

OpenStudy (anonymous):

@sauravshakya

OpenStudy (anonymous):

yea ill try. Im gona look it up

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

This is how i did it, skip down to the sample growth pop problem. If you me to do it though, just ask and i ll try walk you through it.

OpenStudy (anonymous):

So whats the answer ?

OpenStudy (anonymous):

umm 1 sec

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im srry, I only have a graphing calc. not sure how to use it compared to the old ti i had.. :( rly srry. i think the equation is 22=(6.5438x10^4)xe^(.47x20) ? i think. Just trying to help.

OpenStudy (wolf1728):

65,438 population = 65,438 * 1.047^y where y is number of years. so after 1 year = 65,438 * 1.047^1 = 68,514 After 22 years 65,438 * 1.047^22 = 65,438 * 2.7467997357 equals 179,745

OpenStudy (anonymous):

Thank you sooo much . Can you help me with more @wolf1728

OpenStudy (anonymous):

the number of customers visiting a local business is 61 874 and has been continuously declining at a rate of 3.5% each year. What is the approximate number of customers in 15 years?

OpenStudy (anonymous):

@wolf1728